The equation of transverse wave on a rope is
y(x, t) = 5sin (4.0t − 0.02x)
where y and x are measured in cm and t is expressed in second. Calculate the
maximum speed of a particle on the rope
1
Expert's answer
2019-03-04T08:56:57-0500

The wave function for the transverse wave on the rope can be written as follows:


"y(x,t) = A sin(\omega t - kx) = 5 sin(4.0 t - 0.02x) ,"

here, "A = 0.05 m" is the amplitude of the transverse wave, "\omega = 4.0 \dfrac{rad}{s}" is the angular frequency of the transverse wave and "k = 0.02 \dfrac{rad}{cm} \cdot \dfrac{100 cm}{1 m} = 2 \dfrac{rad}{m}" is the wavenumber.

By the definition, the velocity is the derivative of position with respect to time:


"v = \dfrac{dy}{dt}."


So, we can write the speed of a particle on the rope as follows:


"v(x, t) = \dfrac{dy(x,t)}{dt} = \dfrac{d}{dt}(A sin(\omega t - kx)) = A \omega cos(\omega t - kx)."

As we can see from the formula, the speed of a particle on the rope is maximum when "cos(\omega t - kx) = \pm 1".

Therefore,


"v_{max} = A \omega = 0.05 m \cdot 4.0 \dfrac{rad}{s} = 0.2 \dfrac{m}{s}."

Answer:

"v_{max} = 0.2 \dfrac{m}{s}."